Circular No. 7725

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 S1 (SKIFF) B. Skiff, Lowell Observatory, reports his discovery of an apparent comet

on LONEOS telescope images obtained on Sept. 26.4 UT, when the object appeared distinctly less well concentrated than nearby stars (despite poor seeing), with an apparent 17′′ coma. Observations by R. H. McNaught at Siding Spring on Sept. 26 show a coma barely different from the 2′′– 3′′ seeing but with a short tail (< 10′′ long) in p.a. ∼ 40^{o}. J. G. Ries reports that CCD images obtained with the 0.76-m reflector at McDonald Observatory on Sept. 28.4 confirm the northeastward tail of length ∼ 10′′.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Sept.26.37650  2^{h}23^{m}38.^{s}75  −29^{o}48′06′′.3  19.7

All the available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2001-S75:

T = 2001 June 1.104 TT  ω = 284.^{o}221
Ω = 329.579 }2000.0
q = 3.75865 AU  i = 139.348

SUPERNOVA 2001en IN NGC 523 An independent discovery of this supernova (cf. IAUC 7724) at Beijing

Astronomical Observatory has been relayed from Wan Zhou via W. Li. The Beijing group found SN 2001en on an exposure taken on Sept. 27.53 UT at mag ∼ 17.4. Its position end figures were found to be 22^{s}.82, 30′′.2 (equinox 2000.0). Nothing was visible on a Beijing image obtained on Sept. 12.64 (limiting mag 18.2).

V4334 SAGITTARII H. U. Kaeufl, European Southern Observatory (ESO); F. Kerber, Space

Telescope European Coordinating Facility, ESO; and J. Koller, Rice University, report spectrophotometric fluxes, obtained with the ESO 3.6-m telescope (+ new infrared focal reducer TIMMI2): 2000 Oct. 11, 69.1 Jy in band 7.7–8.1 μm, 32.9 Jy in band 11.7–12.9 μm; 2001 Aug. 13, 67.0 Jy, 57.3 Jy. They add: “Compared to the last known measurement in this wavelength regime (1998 June 12; IAUC 6938), the infrared flux density had increased by a factor of seven by 2000 Oct. 11. Thereafter the changes in flux are less drastic and indicate cooling of the circumstellar material.”

2001 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green